OpenStudy (anonymous):

I would like some help on Absolute value equations. Rewards to helpers. I have some example problems here. Thanks! :)

OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12

OpenStudy (youngstudier):

Okay, so what do you think would be the first step?

OpenStudy (anonymous):

Wouldn't we have to find possible solutions?

OpenStudy (youngstudier):

I don't think so.... But we both do different schooling programs.... I'm not entirely sure what your program would say to do for this type of question. Do you know if there are answer choices for the problems or no?

OpenStudy (anonymous):

Yes, there are. One second. Let me pull them up.

OpenStudy (anonymous):

x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (youngstudier):

@campbell_st

OpenStudy (anonymous):

@amistre64

OpenStudy (campbell_st):

the 1st step to solving this is to divide both sides of the equation by -3 this results in \[\left| x + 7 \right| = 4\] when you have the equation to this point, you need to know there will be 2 solutions to this problem the can be found by using x + 7 = 4 and x + 7 = -4 now just solve for x

OpenStudy (anonymous):

Okay, thanks. I'll open up more in a second.